# Patent Language Model Pretraining with ModernBERT

### v2

@@ 3.2 Downstream Evaluation
ModernBERT-base 0.806 0.786 0.773 0.822
ModernBERT-base-PT 0.802 0.814 0.782 0.843
ModernBERT-base-VX 0.796 0.801 0.776 0.852
MosaicBERT-large 0.787 0.796 0.772 0.850
PatentBERT 0.801 0.807 0.810 0.854
Table 3: F1 Scores for downstream classification tasks on four patent datasets.
@@ 3.3 Finetuning Efficiency
ModernBERT has reported efficiencies in downstream training and inference Warner et al. (2024), which are further validated here. Results are reported in Table 4 and plotted with confidence intervals over 4 random seeds in Appendix E. For finetuning tasks, the average number of samples per second processed during training were 33.6, 22.7, 24.0, 9.2 and 11.5 for ModernBERT-base, ModernBERT-base-PT, ModernBERT-base-VX, MosaicBERT-large and PatentBERT, respectively. For inference, these numbers were 154.5, 157.2, 157.1, 53.3 and 44.0 samples per second. These gains stem primarily from the use of FlashAttention, ALiBi positional embeddings, and GLU-based feed-forward layers with results validating the ModernBERT architecture as a possible alternative to other BERT variants, especially in time-critical inference applications.
Model Parameters Training Inference
ModernBERT-base 149M 33.6 154.5
ModernBERT-base-PT 149M 22.7 157.2
ModernBERT-base-VX 149M 24.0 157.1
MosaicBERT-large 340M 9.2 53.3
PatentBERT 346M 11.5 44.0
Table 4: Model parameters and downstream throughput in samples per second.

### v3

@@ 3.2 Downstream Evaluation
ModernBERT-base 0.806 $\pm$ 0.011 0.786 $\pm$ 0.005 0.773 $\pm$ 0.004 0.822 $\pm$ 0.019
ModernBERT-base-PT 0.802 $\pm$ 0.011 0.814 $\pm$ 0.004 0.782 $\pm$ 0.003 0.843 $\pm$ 0.009
ModernBERT-base-VX 0.796 $\pm$ 0.003 0.801 $\pm$ 0.008 0.776 $\pm$ 0.005 0.852 $\pm$ 0.005
MosaicBERT-large 0.787 $\pm$ 0.010 0.796 $\pm$ 0.007 0.772 $\pm$ 0.006 0.850 $\pm$ 0.008
PatentBERT 0.801 $\pm$ 0.013 0.807 $\pm$ 0.006 0.810 $\pm$ 0.001 0.854 $\pm$ 0.012
Table 3: F1-score (mean $\pm$ std across 4 runs) for downstream classification tasks on four patent datasets.
@@ 3.3 Finetuning Efficiency
[…]
Model Parameters WIPO WIPOEC HUPD DatasetCLV
ModernBERT-base-PT 149M 83.68 $\pm$ 0.09 85.54 $\pm$ 0.15 368.42 $\pm$ 0.33 90.99 $\pm$ 0.83
PatentBERT 346M 36.09 $\pm$ 0.06 35.50 $\pm$ 0.12 68.84 $\pm$ 0.38 35.51 $\pm$ 0.05
ModernBERT-base 149M 84.48 $\pm$ 0.04 62.66 $\pm$ 0.52 377.41 $\pm$ 1.73 93.45 $\pm$ 0.42
ModernBERT-base-VX 149M 84.06 $\pm$ 1.04 84.51 $\pm$ 0.38 368.09 $\pm$ 0.75 91.87 $\pm$ 0.23
MosaicBERT-large 340M 27.44 $\pm$ 0.09 26.75 $\pm$ 0.56 127.39 $\pm$ 0.60 31.74 $\pm$ 0.30
Table 4: Inference throughput (samples/sec; mean $\pm$ std over runs) for four patent datasets.
Model Parameters WIPO WIPOEC HUPD DatasetCLV
ModernBERT-base-PT 149M 21.11 $\pm$ 0.43 21.48 $\pm$ 1.01 24.98 $\pm$ 1.58 23.33 $\pm$ 0.16
PatentBERT 346M 8.65 $\pm$ 0.14 9.29 $\pm$ 0.06 19.29 $\pm$ 0.56 8.87 $\pm$ 0.07
ModernBERT-base 149M 21.51 $\pm$ 0.54 23.17 $\pm$ 0.93 65.04 $\pm$ 2.13 24.55 $\pm$ 0.20
ModernBERT-base-VX 149M 20.85 $\pm$ 0.38 21.15 $\pm$ 0.80 30.38 $\pm$ 2.67 23.50 $\pm$ 0.13
MosaicBERT-large 340M 6.19 $\pm$ 0.12 6.53 $\pm$ 0.16 17.33 $\pm$ 0.71 6.92 $\pm$ 0.07
Table 5: Training throughput (samples/sec; mean $\pm$ std over runs) for four patent datasets.
@@ Appendix E Results
Figure 6: Train throughput (samples per second) across models and datasets, reported as mean $\pm$ 95% confidence interval computed over 4 random seeds per (dataset, model). Confidence intervals are estimated using the Student’s $t$ distribution. Higher values indicate greater inference efficiency.